$6,000 dollars is invested in two different accounts earning 3% and 5% interest. At the end of one year, the two accounts earned $220 in interest. How much money was invested at 3%?
--------------------------
it has to be one of the following
--------------------------
$2,000
$3,000
$4,000

Answer by mananth(16946)

Part I 3.00% per annum ------------- Amount invested =x
Part II 5.00% per annum ------------ Amount invested = y
6000
Interest----- 220.00

Part I 3.00% per annum ---x
Part II 5.00% per annum ---y
Total investment
x + 1 y= 6000 -------------1
Interest on both investments
3.00% x + 5.00% y= 220
Multiply by 100
3 x + 5 y= 22000.00 --------2
Multiply (1) by -3
we get
-3 x -3 y= -18000.00
Add this to (2)
0 x 2 y= 4000
divide by 2
y = 2000
Part I 3.00% $ 4000
Part II 5.00% $ 2000

CHECK
4000 --------- 3.00% ------- 120.00
2000 ------------- 5.00% ------- 100.00
Total -------------------- 220.00

m.ananth@hotmail.ca